OpenStudy (anonymous):

how do you graph a equations with the y intercept being 4 and the x intercept -5..how do you know where to start your points

OpenStudy (anonymous):

\[(x _{1}, y _{1}); (x _{2}, y _{2})\]\[m = (y _{1}-y _{2})/(x _{1}-x _{2})\]\[(y-y _{1}) = m(x - x _{1})\]A good way to approach this would be to use point-slope form. Since we know that the y-intercept has an x-value of 0 and that the x-intercept has a y-value of 0, we know that we have 2 points (0, 4) and (-5, 0). Using the slope formula, we find the slope to be (4/5). Now that we have points and slope, we can plug one point and the slope into point-slope form and solve for y: (y - 4) = (4/5)(x - 0) y -4 = (4/5)x y = (4/5)x + 4